OpenStudy (amistre64):

there are 5 questions of a test; each question has 6 choices, of which only 1 is correct. What is the probability of guessing 2 wrongs and 3 correct?

OpenStudy (anonymous):

\[\dbinom{5}{2}(\frac{1}{6})^3\times (\frac{5}{6})^2\]

OpenStudy (anonymous):

good afternoon guru. you wanna number for this or just testing?

OpenStudy (amistre64):

well, im trying to verify my answer to the study guide:) I got 25/7776 for it :)

OpenStudy (amistre64):

how many possible combinations of 2w and 3c are there? that one i get stuck on

OpenStudy (anonymous):

there are \[\dbinom{5}{2}\] of them

OpenStudy (anonymous):

forget choosing 2 and 3. just choose 2 because if you have two you also have the other three. like tossing a coin 5 times and asking how many ways to get 2 heads. same as saying "how many ways to get 2 heads and 3 tails"

OpenStudy (anonymous):

only two possibilities. "bernoulli trials" is the term of art and the formula is \[P(X=k)=\dbinom{n}{k} p^k(1-p)^{n-k}\]

OpenStudy (amistre64):

that looks like the binomial probability distribution :)

OpenStudy (anonymous):

where \[P(X=k)\] means the probability you get exactly k successes in n trials. yes binomial is exactly what you have

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

they don't call me master for nothing!

OpenStudy (amistre64):

lol .... how much do they charge you :)

OpenStudy (anonymous):

plenty and that is just an estimate!

OpenStudy (amistre64):

thanx :) Massa'

OpenStudy (anonymous):

i got \[\frac{125}{2592}\]

OpenStudy (amistre64):

the answer finally emerged later in the day; the first part(a) asked for the probablity of gett 2wrongs and 2 corrects: 1.1.1.5.5 25 -------- = ----- 6^5 7776 the part(b) saisd; how many ways is there to get 2wrong and 3 correct; I kept coming up with 10 and had no clue what to do with it :) cccww ccwcw cwccw wcccw ccwwc cwcwc wccwc cwwcc wcwcc wwccc................................ 10 ways to do it. part(c) what is the probability of getting 2wrong and 3 correct; 25 250 ----- * 10 = ----- ; reduces as wanted 7776 7776 :) thanx

OpenStudy (anonymous):

two wrong and 2 correct? what is the 5th? now answer?

OpenStudy (amistre64):

2wrong and 2 correct out of 5 questions? I dont quite understand the question :)